The son is 23 years younger than the mother. In seven years, the mother will be twice as old as the daughter. How old is each now?
Answer by ankor@dixie-net.com(22740)

The son is 23 years younger than the mother.
In seven years, the mother will be twice as old as the daughter.
How old is each now?
:
Not much information here, let's assume Mommy is 31
then
son is 8 yrs old
and
31 + 7 = 2(d + 7)
31 + 7 = 2d + 14
31 + 7 - 14 = 2d
31 - 7 = 2d
24 = 2d
d = 24/2
d = 12 yrs is the daughter
:
There are many possibilities that are reasonable, and will give an integer value for the daughter's age
Mommy's age could be 27, 29, 31, 33, all odd numbers up to 89 and even beyond
For example when Mom is 89, son will be 66, daughter will be 41
: